solve 6 = -y + 42 - 2y
Answer by Adrish(44)

Solutions:
6 = -y +42 - 2y
6 = -3y +42
6-42 = -3y +42 -42
-36 = -3y
Dividing both sides by -3 , we get
-36/-3 = -3y/-3
12 = y